# Cross Domain Early Crop Mapping using CropGAN and CNN Classifier

### v1

@@ V-A Experiment setup
This work designs two sets of experiments to comprehensively evaluate the cross-domain performance of the proposed CropGAN method. The first set of experiments investigates the cross-year scenario, where Jackson County from 2020 to 2022 is considered the target domain, and Jackson County in 2019 serves as the foundational source domain. In contrast, the second set of experiments focuses on the cross-region scenario, where Cavalier County 2019 and Sussex County 2019 serve as the source domains, respectively. The target domains include the sub-areas from Jilin Province, China, and Quebec, Canada.
In each experiment, the crop mapper is trained by using $70\%$ of randomly selected labelled data from the source domain. The remaining $30\%$ is divided equally into validation and test dataset. To train the CropGAN model, the unlabelled remote sensing data from the target domain is combined with the labelled data from the source domain. During training, the number of epochs is set to $300$, and a batch size of $64$ is used. For optimization, the Adaptive Moment Estimation (Adam) optimizer is employed with a learning rate of $0.005$ and an exponential decay rate of 0.5 for the first moment estimates. It’s worth noting that the coefficients of the total loss function in the experiments are set to $1$, $10$, and $5$.
The results are compared with the direct transfer method as the baseline method, which involves directly applying the trained crop mapper to the target domains without any refinement or preprocessing steps. This comparison enabled us to showcase the effectiveness and improvements achieved by our proposed approach.
V-B Evaluation Metrics
@@
V-C Results
During the training process of CropGAN, a comprehensive evaluation approach is utilized to assess the model’s performance and select the well-trained model.
At the end of each training epoch of CropGAN, data is transformed from the target domain to the source domain. This transformed data is then classified using the crop mapper, and three evaluation metrics, namely the OA, F1 score, and Kappa coefficient, are calculated based on the ground truth.
In addition to the evaluation metrics, we also calculate the total loss between the transformed data and the source domain data to assess how well the CropGAN model aligns the target domain data with the source domain data.
To select the well-trained CropGAN model, the model with the lowest total loss value is chosen. This indicates that the model has successfully minimized the domain shift and effectively transformed the target domain data to resemble the source domain data. The best-trained CropGAN models for the USA target domains (specifically, Jackson County from 2020 to 2022) correspond to epochs 178, 252, and 117. Similarly, the most proficiently trained CropGAN models for the study areas in China and Canada are linked to epochs 238 and 148.
By incorporating both the evaluation metrics and the total loss calculation, our evaluation approach ensures that the selected CropGAN model not only achieves high accuracy in crop mapping but also effectively mitigates the domain shift problem.
TABLE V: The Metrics Comparison of Cross-Year Experiments. Better metrics are indicated in bold.
Source Domain Target Domain CropGAN Baseline
OA F1 Kappa OA F1 Kappa
2019 Jackson 2020 Jackson 0.8149 0.8153 0.6310 0.6162 0.4304 0.2623
2019 Jackson 2021 Jackson 0.8440 0.8347 0.6903 0.5789 0.3249 0.1868
2019 Jackson 2022 Jackson 0.8902 0.8992 0.7792 0.6491 0.4871 0.3141
TABLE VI: The Metrics Comparison of Cross-Region Experiments. ”SA” represents the ”Study Area”. Better metrics are indicated in bold.
Source Domain Target Domain CropGAN Baseline
OA F1 Kappa OA F1 Kappa
2019 Cavalier 2019 SA (China) 0.7155 0.7088 0.4308 0.6268 0.5232 0.2435
2019 Sussex 2019 SA (Canada) 0.7647 0.8174 0.4916 0.7076 0.6861 0.4501
Additionally, experiments are carried out to compare the performance of the direct transfer method and our proposed CropGAN method in crop mapping. For the direct transfer method, the target domain data is directly fed into the crop mapper to provide the crop mapping result. Following this, 3 metrics are calculated as well. The results of the cross-year and cross-region experiments are presented in Table V and VI. The tables provide a comprehensive comparison of various evaluation metrics.
Overall, the CropGAN method consistently outperformed the direct transfer method in terms of 3 metrics. These results highlight the effectiveness of our proposed approach in improving the accuracy of early crop mapping.
In Jackson County in 2020, the CropGAN method achieved a higher OA of 0.8149 compared to the direct transfer method’s OA of 0.6162. Similarly, the CropGAN method exhibited a higher Kappa coefficient value of 0.6310, surpassing the direct transfer method’s Kappa coefficient of 0.2623. The F1 score also showed improvements with the CropGAN method, reaching 0.8153 from 0.4304. For Jackson County in 2021, the CropGAN method outperformed the direct transfer method across all metrics. Notably, the CropGAN method achieved a significantly higher Kappa coefficient of 0.6903 compared to the direct transfer method’s Kappa coefficient of 0.1868. The OA and F1 score also exhibited notable improvements with our proposed approach. In Jackson County in 2022, the CropGAN method outperformed the direct transfer method, reaching a 0.7792 Kappa coefficient compared to 0.3141. Figures 9, 10, and 11 illustrate the visualization of results for the cross-year experiments obtained using both our method and the direct transfer method. In the visualization of mapping results, green pixels represent the identified corn regions, while white pixels represent the non-corn regions. For the error images red highlights the misclassified pixels, while white represents the correctly classified pixels.
Figure 9: Corn Crop Mapping Results Comparison for the Jackson County 2020: Our Method vs. Direct Transfer. (a) displays the ground truth. The crop mapping results are depicted in (b) for CropGAN, and (c) for the baseline method. In this visualization, green denotes corn, and white represents other crops. The corresponding error images are illustrated in panels (d) and (e) for CropGAN and the baseline method, respectively. Red highlights the misclassified pixels, and white represents correctly classified pixels.
Figure 10: Corn Crop Mapping Results Comparison for the Jackson County 2021: Our Method vs. Direct Transfer.
Figure 11: Corn Crop Mapping Results Comparison for the Jackson County 2022: Our Method vs. Direct Transfer.
Moving on to the study area in China, our proposed CropGAN method consistently outperformed the direct transfer method. The improvement is significant, with the CropGAN method achieving an OA of 0.7155 compared to the direct transfer method’s OA of 0.6268. The F1 score and Kappa coefficient also showed remarkable increases with our proposed approach from 0.5232 and 0.2435 to 0.7088 and 0.4308. In the Canadian study area, our proposed CropGAN method consistently outperformed the direct transfer method, with a notable enhancement in the F1 score, which increased from 0.6861 to 0.8174. Figures 12 and 13 illustrate the visualization of results for the cross-domain experiments obtained using both our method and the direct transfer method.
Figure 12: Corn Crop Mapping Results Comparison for the Study Area in China: Our Method vs. Direct Transfer.
Figure 13: Corn Crop Mapping Results Comparison for the Study Area in Canada: Our Method vs. Direct Transfer.
The experimental results demonstrate the superiority of the CropGAN method over the direct transfer method, showcasing significant improvements in OA, F1 score, and Kappa coefficient. By leveraging the combination of labelled data from the source domain and unlabeled data from the target domains, our proposed approach effectively enhances early crop mapping accuracy.
In conclusion, the experimental results validate the effectiveness of our proposed CropGAN method in improving crop mapping accuracy by effectively addressing cross-domain challenges. The consistent outperformance of the CropGAN method compared to the direct transfer method across various metrics highlights its potential for practical applications in the agricultural industry.
V-D The t-SNE Visualization
@@
Figure 14: The t-SNE Visualization of Corn Data Points for the Cross-Year Experiments: Comparison between Target, Transformed Target, and Source Domains. (a) Jackson County 2020. (b) Jackson County 2021. (c) Jackson County 2022.
@@
Figure 15: The t-SNE Visualization of Corn Data Points for the Cross-Region Experiments: Comparison between Target, Transformed Target, and Source Domains. (a) The Study Area in China 2019. (b) The Study Area in Canada 2019.
@@
To assess the effectiveness of our method in addressing the domain shift problem, we utilize t-distributed stochastic neighbour embedding (t-SNE) [31] visualization to analyze the distribution of corn and non-corn data in the target, transformed target, and source domains. By using t-SNE, the data points are projected into a two-dimensional space while preserving their local relationships. Figure 14 and 15 show the t-SNE visualization that illustrates the distribution of corn data points for the cross-domain experiments.
In the visualization, the orange points represent remote sensing data points of corn cropland sourced from the target domain, offering a glimpse into the data distribution within that domain. The green points denote the transformed corn cropland remote sensing data points from the target domain to the source domain, employing our proposed method. Finally, the blue points indicate the original corn cropland remote sensing data points extracted from the source domain.
Upon analyzing the t-SNE visualization, it is evident that the distribution of corn cropland remote sensing data points between the source and target domains differs. This disparity highlights the presence of a domain shift, which poses challenges for accurate crop mapping under the target domains.
@@
However, the application of our CropGAN method resulted in an improvement in the similarity between the data distribution of the transformed target domain data (green points) and the source domain data (blue points), compared to the similarity between the original target domain data (orange points) and the source domain data (blue points). This resemblance enables the crop mapper, trained on the source domain, to effectively process the transformed remote sensing data obtained from the target domain. It coincides with our excellent crop mapping results for these years and this county.
Overall, the t-SNE visualization provides valuable insights into the effectiveness of our method in addressing the domain shift problem and highlights the impact of domain similarity on the performance of crop mapping. Through the utilization of our CropGAN method to bridge the gap between the target and source domains, we attain improved similarity in the distribution between transformed target domain data and source domain data, leading to enhanced early crop mapping accuracy.
VI Discussion and Conclusions
Our study focuses on early crop mapping using a combination of deep learning techniques and domain adaptation. In this paper, a novel approach is introduced that harnesses the capabilities of a CycleGAN-based model, known as CropGAN, to tackle the domain shift problem between source and target domains. By transforming the remote sensing data of the target domain to resemble the source domain, accurate early crop mapping outcomes are achieved.
The experimental results demonstrate the effectiveness of our approach. The performance of our method is compared with a direct transfer method, where the crop mapper is directly applied to the target domain without any preprocessing or refinement. Our method consistently outperformed the direct transfer method across multiple evaluation metrics, including the OA, F1 score, and Kappa coefficient. This indicates that the CropGAN-based domain adaptation significantly improves the accuracy of the early crop mapping in the target domain.
Additionally, the t-SNE visualization results offered insights into the distribution of corn and non-corn data in both the source and target domains. A clear distinction difference between the distributions of the two domains is observed, highlighting the domain shift challenge. However, the distribution of the transformed target domain data by CropGAN closely resembled that of the source domain, signifying the effectiveness of our method in addressing the domain shift problem.
@@ VI Discussion and Conclusions
However, it is important to acknowledge the limitations of our work. One limitation of our CropGAN work is the underlying assumption that the primary crops in the target domain and the source domain are consistent. This assumption, which implies uniformity in the dominant crop types, may not always hold true in practice, particularly when considering diverse agricultural practices across different regions. In future research, addressing this limitation and developing methods that can accommodate variations in primary crop types across domains will be a valuable direction for enhancing the robustness and applicability of our approach.
Overall, our work presents a robust and effective approach to crop mapping by combining deep learning techniques, domain adaptation using CropGAN, and rigorous evaluation metrics. The results demonstrate the superiority of our method compared to direct transfer approaches. This work contributes to the field of crop mapping by addressing the challenges of domain shift and improving the accuracy of crop classification in target domains.

### v2

@@ V Experiment setup and Results
TABLE VI: Cross-Year Experiment Metrics Comparison: Targeting Jackson County 2020. The best metrics are indicated in bold, while the second-best metrics of the baseline methods are underlined.
CropSTGAN CropTGAN STDAN TempCNN
OA F1 Kappa OA F1 Kappa OA F1 Kappa OA F1 Kappa
S1 0.9300 0.9311 0.8596 0.9167 0.9142 0.8339 0.9389 0.9386 0.8769 0.6162 0.4304 0.2623
S2 0.9220 0.9218 0.8441 0.8921 0.8920 0.7843 0.9289 0.9285 0.8565 - - -
S3 0.9266 0.9274 0.8530 0.8861 0.8828 0.7728 0.9248 0.9249 0.8498 - - -
Avg 0.9262 0.9268 0.8522 0.8983 0.8963 0.7970 0.9309 0.9307 0.8611 0.6162 0.4304 0.2623
TABLE VII: Cross-Year Experiment Metrics Comparison: Targeting Jackson County 2021. The best metrics are indicated in bold, while the second-best metrics of the baseline methods are underlined.
CropSTGAN CropTGAN STDAN TempCNN
OA F1 Kappa OA F1 Kappa OA F1 Kappa OA F1 Kappa
S1 0.9212 0.9219 0.8421 0.8884 0.8814 0.7785 0.9199 0.9197 0.8391 0.5789 0.3249 0.1868
S2 0.9182 0.9194 0.8360 0.8837 0.8934 0.7653 0.9006 0.9002 $0$
S3 0.9198 0.9208 0.8393 0.9057 0.8985 0.8124 0.9167 0.9163 0.8322 - - -
Avg 0.9197 0.9207 0.8391 0.8926 0.8911 0.7854 0.9124 0.9121 0.8238 0.5789 0.3249 0.1868
Figure 7: Corn Crop Mapping Results Comparison with Jackson County 2020 as the Target Domain. (a) displays the ground truth. The crop mapping results are depicted in (b) for CropSTGAN, (c) for CropTGAN, (d) for STDAN, and (e) for TempCNN. In this visualization, green denotes corn, and white represents other types. The corresponding error images are illustrated in panels (f) and (i) for CropSTGAN, CropTGAN, STDAN, and TempCNN, respectively. Red highlights the misclassified pixels, and white represents correctly classified pixels.
Figure 8: Corn Crop Mapping Results Comparison with Jackson County 2021 as the Target Domain.
TABLE VIII: Cross-Region Experiment Metrics Comparison: Targeting China 2019. The best metrics are indicated in bold, while the second-best metrics of the baseline methods are underlined.
CropSTGAN CropTGAN STDAN TempCNN
OA F1 Kappa OA F1 Kappa OA F1 Kappa OA F1 Kappa
S1 0.8694 0.8762 0.7401 0.8228 0.7961 0.6483 0.7839 0.7289 0.5622 0.6268 0.5232 0.2435
S2 0.8447 0.8239 0.6874 0.7933 0.7533 0.5903 0.7480 0.6732 0.4888 - - -
S3 0.8331 0.8061 0.6634 0.7971 0.7578 0.5978 0.7090 0.6084 0.4084 - - -
Avg 0.8491 0.8354 0.6970 0.8044 0.7691 0.6121 0.7470 0.6702 0.4865 0.6268 0.5232 0.2435
TABLE IX: Cross-Region Experiment Metrics Comparison: Targeting Jackson County 2019. The best metrics are indicated in bold, while the second-best metrics of the baseline methods are underlined.
CropSTGAN CropTGAN STDAN TempCNN
OA F1 Kappa OA F1 Kappa OA F1 Kappa OA F1 Kappa
S1 0.8303 0.8181 0.6630 0.7881 0.7899 0.5815 0.7655 0.7651 0.5286 0.6662 0.6905 0.3282
S2 0.8303 0.8188 0.6627 0.7806 0.8013 0.5569 0.7568 0.7744 0.5106 - - -
S3 0.8190 0.8330 0.6347 0.7914 0.7970 0.5821 0.7216 0.6883 0.4534 - - -
Avg 0.8265 0.8233 0.6535 0.7867 0.7961 0.5735 0.7480 0.7426 0.4975 0.6662 0.6905 0.3282
Figure 9: Corn Crop Mapping Results Comparison with the Study Area of China 2019 as the Target Domain.
Figure 10: Corn Crop Mapping Results Comparison with Jackson County 2019 as the Target Domain.
@@ V-A Experiment Setup
This work designs two sets of experiments to comprehensively evaluate the cross-domain performance of the proposed CropSTGAN method:
  • The first set of experiments investigates the cross-year scenario, where Jackson County from 2020 to 2021 is considered as the target domain, and Jackson County in 2019 serves as the source domain. The discrepancies in data distribution between these domains are relatively slight.
  • The second series of experiments explores cross-region scenarios, incorporating a study area in China and Jackson County. When Jackson County 2019 serves as the source domain, the target domain is the study area in China 2019. Conversely, when the study area in China 2019 is the source domain, Jackson County 2019 becomes the target domain. The discrepancies in data distribution across the domains are relatively large.
To evaluate the distinct architecture of the CropSTGAN domain mapper, comparative experiments were undertaken utilizing a simpler, analogous structure named CropTGAN. Within this framework, both the discriminators’ and generators’ encoders share the same filter numbers and sizes as those in CropSTGAN but focus solely on extracting features from the temporal dimension. Additionally, we benchmarked the CropSTGAN framework against various SOTA methods, including TempCNN and STDAN.
V-B Training Setup
In each experiment, $200,000$ data points are randomly sampled from both the source and target domains, resulting in a total of $400,000$ data points. The TempCNN is trained using $70\%$ of randomly selected labelled data from the source domain. The remaining $30\%$ is divided equally into validation and test datasets. To train the CropSTGAN model, all the sampled data from the target domain and the source domain are used. During training, a batch size of $64$ is used. All Methods’ training persisted until the completion of 500 epochs or upon convergence, as determined by an early stopping criterion set at 50 epochs. For optimization, the Adaptive Moment Estimation (Adam) optimizer is employed with a learning rate of $0.005$ and an exponential decay rate of $0.9$ for the first moment estimates.
Every method was repeatedly trained on each subset from scratch 3 times with the same training configuration. Notably, the TempCNN model was only trained once and served as the crop mapper for the CropSTGAN, resulting in a single TempCNN outcome for each test. Moreover, in the CropSTGAN and CropTGAN framework training processes, it’s worth noting that the coefficients of the total loss function in the first set of experiments are set to $\alpha=1,\beta=10,\sigma=5$. For the second set of experiments, they are set to $\alpha=1,\beta=20,\sigma=10$. The choice of coefficients in our total loss function is based on practical experience and is designed to ensure the stability and effectiveness of the CropSTGAN and CropTGAN training processes.
V-C Evaluation Metrics
@@
The Kappa coefficient
V-D Results
@@
Figure 11: The t-SNE Visualization of Corn Data Points for the Cross-Year Experiments: Comparison between Target Domain Data, Transformed Target Domain Data, and Source Domain Data. (a) Jackson County 2020 as the target domain. (b) Jackson County 2021 as the target domain.
@@
Figure 12: The t-SNE Visualization of Corn Data Points for the Cross-Domain Experiments: Comparison between Target Domain Data, Transformed Target Domain Data, and Source Domain Data. (a) The study area of China as the target domain. (b) Jackson County 2019 as the target domain.
Tables VI and VII present the results for the first set of cross-year experiments. Tables VIII and IX present the results for the second set of cross-region experiments. Across two sets of experiments, our CropSTGAN method demonstrates the highest average metrics, except for the cross-year experiment targeting Jackson County 2020.
In the first set of experiments targeting Jackson County 2020, the CropSTGAN achieved remarkable results, securing second place with an average OA of 92.62%, an F1 score of 92.68%, and a Kappa coefficient of 85.22%. These results represent significant improvements over TempCNN, with increases of +31.00% in OA, +49.64% in F1, and +58.99% in Kappa. Additionally, CropSTGAN outperformed CropTGAN, underscoring the effectiveness of its unique structure. However, it was slightly outpaced by the STDAN, which took first place with an OA of 93.09%, an F1 score of 93.07%, and a Kappa of 86.11%. When Jackson County 2021 plays as the target domain, CropSTGAN outperforms STDAN, leading by a slight margin with an increase of +0.73% in OA, +0.86% in F1 Score, and +1.53% in Kappa Coefficient. Meanwhile, CropTGAN ranks third, significantly surpassing TempCNN in performance. Figures 7 and 8 display the visualization of cross-year experimental results and their error images using CropSTGAN, CropTGAN, STDAN, and TempCNN.
In the second set of experiments, CropSTGAN outperformed all, securing the top spot, followed by CropTGAN in both tests. Specifically, when targeting the study area of China, CropSTGAN achieved an OA of 84.91%, an F1 score of 83.54%, and a Kappa coefficient of 69.70%. These metrics are markedly higher than those of STDAN, showing increases of 10.21% in OA, 16.52% in F1, and 21.05% in Kappa. Conversely, with China as the source domain and Jackson County 2019 as the target domain, CropSTGAN reached an OA of 82.65%, an F1 score of 82.33%, and a Kappa coefficient of 65.35%, surpassing STDAN by 7.85% in OA, 8.07% in F1, and 15.6% in Kappa. Additionally, CropTGAN, CropGAN, and STDAN all performed better than TempCNN, which only managed an OA of 66.62%, an F1 score of 69.05%, and a Kappa coefficient of 32.82%. Figures 9 and 10 present the results and error images from cross-region experiments.
V-E The t-SNE Visualization
@@
To assess the effectiveness of our CropSTGAN method in addressing the domain shift problem, t-distributed stochastic neighbour embedding (t-SNE) [30] is utilized for the visualization to analyze the distribution of the target data, transformed target data, and source domain data of the target crop. By using t-SNE, the data points are projected into a two-dimensional space while preserving their local relationships. Figure 11 and 12 show the t-SNE visualization that illustrates the distribution of corn data points for the cross-year and cross-region experiments.
@@
In the visualization, the orange points represent MSI data points of corn cropland sourced from the target domain, offering a glimpse into the data distribution within that domain. The green points denote the transformed corn cropland MSI data points from the target domain to the source domain, employing our proposed method. Finally, the blue points indicate the original corn cropland MSI data points extracted from the source domain. Upon analyzing the t-SNE visualization, it is evident that the distribution of corn cropland MSI data points between the source and target domains differs. This disparity highlights the presence of a domain shift, which poses challenges for accurate crop mapping under the target domains. However, the application of our CropSTGAN domain mapper resulted in an improvement in the similarity between the data distribution of the transformed target domain data (green points) and the source domain data (blue points), compared to the similarity between the original target domain data (orange points) and the source domain data (blue points). This resemblance enables the TempCNN crop mapper, trained on the source domain, to effectively process the transformed remote sensing data obtained from the target domain. It coincides with our excellent crop mapping results for these years and this county.
VI Discussion
VI-A Analysis of crop mapping results
In the cross-year experiments, CropSTGAN and STGAN showed similar performance, indicating that CropSTGAN effectively addresses the inter-annual cross-domain challenge, comparable to the state-of-the-art (SOTA) method. In cross-region experiments, CropSTGAN achieved better results than STGAN. As evident from Figure 2, the NDVI discrepancy is more pronounced in cross-region experiments than in cross-year ones. Under the significant differences in data distribution, CropSTGAN outperforms STGAN, benefiting from the identity loss.
Moreover, CropSTGAN outperformed CropTGAN for all experiments, suggesting the effectiveness of the CropSTGAN domain mapper structure. Additionally, the results from cross-domain methods were significantly better than those obtained by directly applying a CNN-based crop classifier, trained on the source domain, to the target domain.
VI-B Advantages of CropSTGAN
Most existing crop mapping related studies rely heavily on a large number of local labelled data for modeling and making predictions, and thus tedious and costly sample collection needs to be carried out extensively and frequently. The sharing of collected labelled samples is an effective way to address the dilemma of ground truth sampling. However, due to the differences in climate conditions across regions and years, trained crop classification models may lose their validity when applied to new domains. Therefore, the CropSTGAN was developed to address the distribution discrepancy existed between the source and target domains, that is, cross-domain issue.
In order to address this issue, most SOTA methods, like STDAN, strive to extract invariant features across target and source domains to tackle cross-domain challenges. However, their effectiveness is often limited by significant differences in data distribution between these domains. In contrast, our CropSTGAN demonstrates superior performance in scenarios with large data distribution disparities, as evidenced by our experimental results.
VI-C Limitations of CropSTGAN
@@
However, it is important to acknowledge the limitations of our work. One limitation of our CropSTGAN work is the underlying assumption that the primary crops in the target domain and the source domain are consistent. This assumption, which implies uniformity in the dominant crop types, may not always hold true in practice, particularly when considering diverse agricultural practices across different regions. In future research, addressing this limitation and developing methods that can accommodate variations in primary crop types across domains will be a valuable direction for enhancing the robustness and applicability of our approach.
VII Conclusions
In conclusion, we introduced the CropSTGAN framework, which integrates a pre-processor, a domain mapper, and a TempCNN crop mapper, specifically designed to overcome the challenges of cross-domain early crop mapping caused by inter-region and inter-year variations. Notably, the framework is versatile, enabling not only crop mapping but also the classification of various land cover types. The CropSTGAN domain mapper is designed to extract both temporal and spectral features from time-series MSI data, effectively transforming target domain data to the source domain. The TempCNN crop mapper, trained by the labelled source domain data, takes the transformed target domain data as input to locate the target crop for the target domain.
Our comprehensive evaluation, conducted across various regions in the USA and China and spanning different years, demonstrates the CropSTGAN framework’s superior performance. It outperforms several baseline and SOTA methods, including TempCNN and STDAN, thereby validating its effectiveness and accuracy in cross-domain early crop mapping scenarios, even with large data distribution disparities between the target domain and source domain.